Solve for x by completing the square: fx^2+gx+h=0
Answer by stanbon(75887)

Solve for x by completing the square: fx^2+gx+h=0
-----
fx^2 + gx = -h
f(x^2 + (g/f)x + ?) = -h + f*?
Complete the square:
----
f(x^2 + (g/f)x + (g/(2f))^2) = -h + f(g/(2f))^2
---
Factor:
f(x+(g/(2f))^2 = -h + (fg^2)/(4f^2)]
----
(x+(g/(2f))^2 = (-h/f) + (g^2)/(4f^2)
Now solve for "x"
================================================
cheers,
Stan H.